Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Interesting GMATFix Problem-51

Expert replies
by arora007 » Sat Oct 09, 2010 2:04 am
On Monday, the probability of rain in NY is 20% and is Paris is 50% and in Moscow is 80%. If these probabilities are independent, what is the probability that it rains on monday at exactly 2 of these 3 locations?
A)1/50
B)8/25
C)2/5
D)21/50
E)1/2
https://www.skiponemeal.org/
https://twitter.com/skiponemeal
Few things are impossible to diligence & skill.Great works are performed not by strength,but by perseverance

pm me if you find junk/spam/abusive language, Lets keep our community clean!!
Join the discussion
Source: — Problem Solving |

by sanju09 » Sat Oct 09, 2010 2:22 am
arora007 wrote:On Monday, the probability of rain in NY is 20% and is Paris is 50% and in Moscow is 80%. If these probabilities are independent, what is the probability that it rains on monday at exactly 2 of these 3 locations?
A)1/50
B)8/25
C)2/5
D)21/50
E)1/2

There are only three ways to have rain at exactly 2 of these 3 locations, and those are:

• NY, Paris rain and Moscow doesn’t, probability is = 0.2 × 0.5 × 0.2 = 0.02

• NY, Moscow rain and Paris doesn’t, probability is = 0.2 × 0.8 × 0.5 = 0.08

• Moscow, Paris rain and NY doesn’t, probability is = 0.8 × 0.5 × 0.8 = 0.32


Total probability = 0.02 + 0.08 + 0.32 = 0.42 = [spoiler]21/50


D, and this is my 2500th post to the forum.:wink:
[/spoiler]
The mind is everything. What you think you become. –Lord Buddha



Sanjeev K Saxena
Quantitative Instructor
The Princeton Review - Manya Abroad
Lucknow-226001

www.manyagroup.com
Join the discussion

by arora007 » Sat Oct 09, 2010 11:16 am
Congrats Sanju...
I guess I am lucky to have ur 2500 post on my post...
I hope to see a 51 in my Q :)
https://www.skiponemeal.org/
https://twitter.com/skiponemeal
Few things are impossible to diligence & skill.Great works are performed not by strength,but by perseverance

pm me if you find junk/spam/abusive language, Lets keep our community clean!!
Join the discussion

by Arcane66 » Sat Oct 09, 2010 12:51 pm
I also got D but it took a bit of thinking. I separated it into three possibilities. Rain in: 1. NY & P but not M. 2. NY & M but not P. 3. M & P but not NY. If you multiply by using the probability that the first two happen times the probability that the third rain doesn't happen, then you will get 21/50 once you total up the probabilities.
Join the discussion